OpenStudy (ebonyserrano):

A light bulb consumes 2400 watt-hours per day. How long does it take to consume 13200 watt-hours? _days _hours

OpenStudy (benji):

Alright first we need to think how many hours are in a day and how many watt-hours are consumed per hour. How would you do that?

OpenStudy (campbell_st):

well if it consumes 2400 watt hours in a day \[2400\div 24 = watts/hour\] then divide this number into 13200 and that will give the number of hours convert it to days and hours an alternative is \[13200 \div 2400 = watt/day\] if you get a decimal part, multiply the decimal by 24 e.g.... 4.25 days = 4 days + 0.25 x 24 or 4 days 6 hours

OpenStudy (danjs):

For the constant rate of power consumption 2400 watt hours per day is the same as 13200 watt hours per how many days \[\large \frac{ 2400~~W*hr }{ 1~~day }=\frac{ 13200~~W*hr }{ X~~day }\]